# Risk-Calibrated Human-Robot Interaction via Set-Valued Intent Prediction

### v1

@@ VII Conclusion
We propose Risk-Calibrated Interactive Planning (RCIP), a framework that applies statistical multi-hypothesis risk control to address the problem of risk calibration for interactive robot tasks. We formalize RCIP as providing a statistical guarantee on an arbitrary number of user-specified risks, such as prediction failures and the amount of human help, subject to a bound on the rate at which the robot fails to predict the optimal actions. By optimizing preferences over a small number of model parameters, RCIP is able to achieve higher flexibility in aligning to user preferences than fixed-paramter methods. Experiments across a variety of simulated and hardware setups demonstrate that RCIP does not exceed user-specified risk levels. Moreover, RCIP reduces user help $5-30\%$ when compared to baseline approaches that lack formal assurances.

### v2

@@ VII Conclusion
We propose Risk-Calibrated Interactive Planning (RCIP), a framework that applies statistical multi-hypothesis risk control to address the problem of risk calibration for interactive robot tasks. We formalize RCIP as providing a statistical guarantee on an arbitrary number of user-specified risks, such as prediction failures and the amount of human help, subject to a bound on the rate at which the robot fails to predict the optimal actions. By optimizing preferences over a small number of model parameters, RCIP is able to achieve higher flexibility in aligning to user preferences than fixed-paramter methods. Experiments across a variety of simulated and hardware setups demonstrate that RCIP does not exceed user-specified risk levels. Moreover, RCIP reduces user help $8-87\%$ when compared to baseline approaches that lack formal assurances.